OpenStudy (anonymous):

helpp!!! please ASAP. the question is : Write an equation for the translation of y=4/x that has the given asymptote: x=1 and y=2

OpenStudy (solomonzelman):

when you graph y=4/x you see asymptotes y=0 and x=0. All you need to do, is to shift the graph 2 units up and one unit right.

OpenStudy (anonymous):

thank you this help but i want to wrtie an equation for the translation

OpenStudy (solomonzelman):

1) shifting f(x) to the left by c units. (x+c) 2) shifting f(x) to the right by c units. (x-c) 3) shifting f(x) c units up. add c to the function. 3) shifting f(x) c units down subtract c to the function.

OpenStudy (solomonzelman):

\(\LARGE\color{black}{ y=\frac{4}{x} }\) to shift it c units up. \(\LARGE\color{black}{ y=\frac{4}{x} +C}\) to shift it c units down. \(\LARGE\color{black}{ y=\frac{4}{x} -C}\) to shift it c units right. \(\LARGE\color{black}{ y=\frac{4}{x-C} }\) to shift it c units left. \(\LARGE\color{black}{ y=\frac{4}{x+C} }\)

OpenStudy (solomonzelman):

You need to do 2 shifts. 1) 1 unit to the right. 2) 2 units up.

OpenStudy (anonymous):

okay thanks this really helped alot thanks

OpenStudy (solomonzelman):

Anytime:) Can you tell me your answer, so that I can verify that you got it? (Please)

OpenStudy (anonymous):

okay hold up

OpenStudy (solomonzelman):

you can draw it, or use Latex.

OpenStudy (solomonzelman):

or screenshot

OpenStudy (anonymous):

Y=(4/x+1) +2

OpenStudy (anonymous):

\[y=(4/x+1)+2\]

OpenStudy (solomonzelman):

Yes

OpenStudy (solomonzelman):

https://www.desmos.com/calculator/r8cuhqijf0

OpenStudy (anonymous):

thank you so muchhh

OpenStudy (solomonzelman):

Yes... and desmos.com is a good graphing calculator for visualizing your problems. (If you need some help graphing something, and I am on, msg me about it.)

OpenStudy (solomonzelman):

Piece wise functions are harder to graph.

OpenStudy (anonymous):

okay thanks i wil when wver i need more help

OpenStudy (solomonzelman):

You are a good learner, I enjoyed teaching even though I myself am a student.... see you.....